A car Is rounding a Circular road of radius 50m at a constant speed of 25 m/s what is the magnitude of the acceleration and it’s acceleration ?
1
Expert's answer
2019-12-18T11:29:38-0500

The centripetal acceleration


"a=\frac{v^2}{R}=\frac{(25\:\rm m/s)^2}{50\:\rm m}=12.5 \:\rm m/s^2"